OpenStudy (anonymous):

Alright guys Trivial question! Find the Definite Integral of Sin^7x dx

OpenStudy (anonymous):

Indefinite*

OpenStudy (anonymous):

doesn't look trivial to me...

OpenStudy (lgbasallote):

it's just a lot of trig transformations =_=

OpenStudy (anonymous):

Hey, it's fun:)

OpenStudy (anonymous):

be my guest... without using Mr. Wolfram...

OpenStudy (anonymous):

Integration by parts.

OpenStudy (anonymous):

Here, take u = sin^6 x and dv= sinxdx

OpenStudy (anonymous):

\[du= 6\sin^5xcos x dx\]

OpenStudy (anonymous):

v= -cosx

OpenStudy (anonymous):

so you're integral = \[-cosxsin^6 x - \int\limits_{}^{} -6cosxsin^5xcosxdx\]

OpenStudy (anonymous):

\[= -\sin^6xcosx + 6\int\limits_{}^{} \sin^5x(\cos^2x)dx\]

OpenStudy (anonymous):

\[= -\sin^6xcosx +6\int\limits_{}^{} \sin^5x (1-\sin^2x) dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\sin^7dx= -\sin^6xcosx + 6\int\limits_{}^{} \sin^5x - \sin^7x dx\]

OpenStudy (anonymous):

Break up the two integrals on the right hand side of the equal sign

OpenStudy (anonymous):

\[ \int\limits_{}^{} \sin^7xdx = -\sin^6xcosx + 6\int\limits_{}^{} \sin^5xdx - 6\int\limits_{}^{} \sin^7x dx\]

OpenStudy (anonymous):

Notice we have two like terms? Put them together to get

OpenStudy (anonymous):

\[7\int\limits_{}^{}\sin^7dx = -\sin^6xcosx+ 6\int\limits_{}^{}\sin^5x\]

OpenStudy (anonymous):

Divide everything by seven to get

OpenStudy (anonymous):

\[\int\limits_{]}^{}\sin^7x dx= -1/7\sin^6xcosx+6/7\int\limits_{}^{}\sin^5x\]

OpenStudy (anonymous):

And that's your answer. Trivia!!

OpenStudy (anonymous):

I don't even know who Mr. Wolfram is.

OpenStudy (anonymous):

i wonder if u can do this with the tabular method to continue on to do that other integral?

OpenStudy (anonymous):

I know that form is accepted, and you don't have to go on finding the integral. But in som it may be easy so one can do it.

OpenStudy (anonymous):

good work man...:)

OpenStudy (anonymous):

Haha thanks, it's quite long I must say.

OpenStudy (anonymous):

YA and i guess it's even longer when you use substitution method

OpenStudy (anonymous):

Definitely wouldn't want to try that, but there's guidelines for when to use trig substitutions

OpenStudy (anonymous):

can u tell me what is that?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

When dealing with trig integrals, If: The power of cosine is odd: Save a cosine and change everything else to sine and then take u=sin x The power of sine is odd: Save a sine and change everything else to cosine then take u=cosx The power of cosx and sinx are both even: Use half-angle formulas. The power of taangent is odd: In most cases, useful substitution is u= secx The power of secant is even: Save a sec^2 x and change everything else to tangent. Take u= tan x. Those are the guidelines

OpenStudy (anonymous):

sin^7x dx = sin^6(x) sin(x)dx = (sin^2x)^3 sinx dx = (1-cos^2x) ^3 sinx dx let u = cosx du = -sinx dx therefore, we get: (1 - u^2) ^3 du = 1 - 3u^2 + 3u^4 - u^5 du integrate from there.

OpenStudy (anonymous):

thanks beeqay

OpenStudy (anonymous):

actually i usually get confused in such type of stuffs

OpenStudy (anonymous):

no problem and @Hermeezey you forgot that it would be -du

OpenStudy (anonymous):

oh yea sorry, I'm easily confused when I'm typing numbers on the computer

OpenStudy (anonymous):

Meant to ask for the solution using integration by parts, because it's more challenging.